A radio tower which is 40 meters in langth is braced with a cable and is tied 9 meters from the base.Find the length of the cable.(Discarded the length of the cable that will be needed for trying.)
Answer by checkley77(12844)

A^2+B^2=C^2
40^2+9^2=C^2
1,600+81=C^2
1,681=C^2
C=SQRT1,681
C=41 METERS IS THE LENGTH OF THE CABLE.